#include <stdio.h>

void printHelp () {
  printf ("\n");
  printf ("a: a(x) = x*x\n");
  printf ("b: b(x) = x*x*x\n");
  printf ("c: c(x) = x^2 + 2*x + 7\n");
  printf ("q: quit\n");
}

void a(float x) {
  float v = x*x;
  printf ("  a(%.2f) = %.2f^2 = %.2f\n", x, x, v);
} // end function a

void b(float x) {
  float v = x*x*x;
  printf ("  a(%.2f) = %.2f^3 = %.2f\n", x, x, v);
} // end function b

void c(float x) {
  float v = x*x + 2*x + 7;
  printf ("  c(%.2f) = %.2f^2 + 2*%.2f + 7 = %.2f\n",
            x, x, x, v);
} // end function c

int menu () {
  char selection;
  float x;
  printHelp ();
  scanf ("%s", &selection);
  if (selection == 'q') return 1;
  scanf ("%f", &x);
  if (selection == 'a') a(x);
  if (selection == 'b') b(x);
  if (selection == 'c') c(x);
  return 0;
} // end function menu

int main() {
  while (menu() == 0);
  printf ("... bye ...\n");
  return 0;
} // end main

Why is the number of x’s different in the printf statements in the functions?
Answer #1
void a(float x) {
  float v = x*x;
  printf ("  a(%.2f) = %.2f^2 = %.2f\n", x, x, v); // it is equal to a(x) = x^2 = v(=x*x)
} // end function a

In the above printf statement, x value is used two times and v value is used one time. 
The expression inside printf statement prints value of function a(x).
The expression is : a(x)=x^2=v
Hence, two times x value and one time v value is passed.


void b(float x) {
  float v = x*x*x;
  printf ("  b(%.2f) = %.2f^3 = %.2f\n", x, x, v); // it is equal to b(x) = x^3 = v(=x*x*x)
} // end function b

In the above printf statement, x value is used two times and v value is used one time.
The expression inside printf statement prints value of function b(x).
The expression is : b(x)=x^3=v
Hence, two times x value and one time v value is passed.


void c(float x) {
  float v = x*x + 2*x + 7;
  printf ("  c(%.2f) = %.2f^2 + 2*%.2f + 7 = %.2f\n",
            x, x, x, v); // it is equal to c(x) = x^2+2*x+7 = v(=x*x + 2*x + 7)
} // end function c 

In the above printf statement, x value is used three times and v value is used one time.
The expression inside printf statement prints value of function c(x).

The expression is : c(x)=x^2+2*x+7

Hence, three times x value and one time v value is passed.

To better understand , count the number of occurrences of x and v in the expression.